### v1

# Dynamic Pricing with Adversarially-Censored Demands

Abstract
We study an online dynamic pricing problem where the potential demand at each time period $t=1,2,\ldots,T$ is stochastic and dependent on the price. However, a perishable inventory is imposed at the beginning of each time $t$, censoring the potential demand if it exceeds the inventory level. To address this problem, we introduce a pricing algorithm based on the optimistic estimates of derivatives. We show that our algorithm achieves $\tilde{O}(\sqrt{T})$ optimal regret even with adversarial inventory series. Our findings advance the state-of-the-art in online decision-making problems with censored feedback, offering a theoretically optimal solution against adversarial observations.
1 Introduction
The problem of dynamic pricing, where the seller proposes and adjusts their prices over time, has been studied since the seminal work of Cournot, (1897). The crux to pricing is to balance the profit of sales per unit with the quantity of sales. Therefore, it is imperative for the seller to learn customers’ demand as a function of price (commonly known as the demand curve) on the fly. However, the demand can often be obfuscated by the observed quantity of sales, especially when censored by inventory stockouts. Such instances severely impede the seller from learning the underlying demand distributions, thereby hindering our pursuit of the optimal price.
Existing literature has devoted considerable effort to the intersection of pricing and inventory decisions. Such works often consider scenarios with indirectly observable lost demands (Keskin et al.,, 2022), recoverable leftover demands Chen et al., (2019), or controllable inventory level (Chen et al.,, 2024). However,these assumptions do not always align with the realities faced in various common business environments. To illustrate, we present two pertinent examples:
Example 1.1 (Performance Tickets).
Imagine that we manage a touring company that arranges a series of performances featuring a renowned artist across various cities. Each venue has a different seating capacity, which substantially affects how we set ticket prices. If the price is too high, it might deter attendance, leading to lower revenue. On the other hand, setting it too low could mean that tickets sell out quickly, leaving many potential attendees unable to purchase them. We do not know exactly how many people attempt to buy tickets and fail. Moreover, because the performances are unique, there is no assurance that those who miss out on one show will choose or be able to attend another. This variability in venue size across different locations requires us to continually adapt our pricing strategy, ensuring that we maximize both attendance and revenue while accommodating unpredictable changes in seat availability.
Example 1.2 (Fruit Retails).
Sweetsop (Annona squamosa, or so-called "sugar apple") is a particularly-perishable tropical fruit, typically lasting only 2 to 4 days (Crane et al.,, 2005). Suppose we manage a local fruit shop and have partnered with a nearby farm for the supply of sweetsops during the harvest season. Due to the perishability nature, we receive sweetsops as soon as they are ripen and picked from the farm every day. This irregular supply means that some days we might receive a large quantity while getting very few on other days. We must quickly sell these fruits before they spoil, yet managing the price becomes challenging. If we exhaust our inventory ahead of time, customers will turn to other fruit shops to make their purchases rather than waiting for our next restock.
Products in the two instances above have the following properties,
  1. 1.
    Inventory level is determined by natural factors, and is arbitrarily given for every individual time period.
  2. 2.
    Products are perishable and salable only within a single time period.
In this work, we study a dynamic pricing problem where the products possess these properties. The problem model is defined as follows. At each time $t=1,2,\ldots,T$, we firstly propose a price $p_{t}$, and then a price-dependent potential demand occurs as $d_{t}$. However, we might have no access to $d_{t}$ as it is censored by an adversarial inventory level $\gamma_{t}$. Instead, we observe a censored demand $D_{t}=\min\{\gamma_{t},d_{t}\}$ and receive the revenue $r_{t}$ as a reward at $t$. Our goal is to approach the optimal price $p_{t}^{*}$ at every time $t$, thereby maximizing the cumulative revenue.
Dynamic pricing with adversarial inventory constraint. For $t=1,2,...,T:$ 1. The seller (we) receives $\gamma_{t}$ identical products. 2. The seller proposes a price $p_{t}\geq 0$. 3. The customers generate an invisible potential demand $d_{t}\geq 0$, dependent on $p_{t}$. 4. The market reveals an inventory-censored demand $D_{t}=\min\{\gamma_{t},d_{t}\}$. 5. The seller gets a reward $r_{t}=p_{t}\cdot D_{t}$. 6. All unsold products perish before $t+1$.
1.1 Summary of Contributions
We consider the problem setting displayed above and assume the potential demand $d_{t}=a-bp_{t}+N_{t}$ is linear and noisy. Here $a,b\in\mathbb{R}^{+}$ are fixed unknown parameters and $N_{t}$ is an unknown i.i.d. [^1] Independently and identically distributed noise with zero mean. Under this premise, the key to deriving the optimal price is to accurately learn the expected reward function $r(p)$, which is equivalent to learning the linear parameters $[a,b]$ and the noise distribution. We are confronted by three principal challenges:
  1. 1.
    The absence of unbiased observations of the potential demand or its derivatives with respect to $p$, which prevents us from estimating $[a,b]$ directly.
  2. 2.
    The dependence of the optimal prices on the noise distribution, which is assumed to be unknown and partially censored.
  3. 3.
    The arbitrariness of the inventory levels, leading to non-stationary and highly-differentiated optimal prices $\{p_{t}^{*}\}$ over time.
[…]
Novelty.
To the best of our knowledge, we are the first to study the online dynamic pricing problem under adversarial inventory levels. Our C20CB algorithm attains an optimal $\tilde{O}(\sqrt{T})$ regret guarantee with high probability. The methodologies we develop are not only pivotal to our design and analysis but also potential for broad application in a variety of online decision-making scenarios with censored feedback.
1.2 Paper Structure
The rest of this paper is organized as follows. We discuss and compare with related works in Section 2, and then describe the problem setting in Section 3. We propose our main algorithm C20CB in Section 4 and analyze its regret guarantee in Section 5. We further discuss the limitations and potential extensions of our methodologies in Section 6, followed by a brief conclusion in Section 7.
2 Related Works
There exists a large volume of literature related to the problem we study in this work. Here we discuss them in the following categories.
Data-driven dynamic pricing
Dynamic pricing for identical products is a well-established research area, starting with Kleinberg and Leighton, (2003) and continuing through seminal works by Besbes and Zeevi, (2009); Broder and Rusmevichientong, (2012); Wang et al., (2014, 2021). The standard approach involves learning a demand curve from price-sensitive demand arriving in real-time, aiming to approximate the optimal price. Kleinberg and Leighton, (2003) provided algorithms with regret bounds of $O(T^{\frac{2}{3}})$ and $O(\sqrt{T})$ for arbitrary and infinitely smooth demand curves, respectively. Wang et al., (2021) refined this further, offering an $O(T^{\frac{k+1}{2k+1}})$ regret for $k$-times continuously differentiable demand curves. This line of inquiry is also intricately linked to the multi-armed bandit problems (Lai and Robbins,, 1985; Auer et al.,, 2002) and continuum-armed bandits (Kleinberg,, 2004), where each action taken reveals a reward without insight into the foregone rewards of other actions.
Contextual pricing: Linear valuation and binary-censored demand
A surge of research has delved into feature-based dynamic pricing (Cohen et al.,, 2020) or pricing with contexts/covariates (Amin et al.,, 2014; Miao et al.,, 2019; Liu et al.,, 2021). These works considered situations where each pricing period was preceded by a context, influencing both the demand curve and noise distribution. Specifically, Cohen et al., (2020); Javanmard and Nazerzadeh, (2019); Xu and Wang, (2021) explored a linear valuation framework with known distribution noise, leading to binary customer demand outcomes based on price comparisons to their valuations. Expanding on this, Golrezaei et al., (2019); Fan et al., (2024); Luo et al., (2021); Xu and Wang, (2022) examined similar models but with unknown noise distributions. In another vein, Ban and Keskin, (2021); Wang et al., (2025); Xu and Wang, (2024) investigated personalized pricing where demand was modeled as a generalized linear function sensitive to contextual price elasticity. Many of these works on valuation-based contextual pricing also assume a censored demand: The seller only observes a binary feedback determined by a comparison of price with valuation, instead of observing the valuation directly. However, it was important to differentiate between the linear (potential) demand model we assumed and their linear valuation models, and there exists no inclusive relationship from each other.
Network Revenue Management (NRM)
NRM (Talluri and Van Ryzin,, 2006) studies the pricing and allocation problem on shared resources in a network. In the problem settings of Besbes and Zeevi, (2012) and the following works Simchi-Levi et al., (2019); Miao and Wang, (2024), if we take a marginal observation from each different product, the occupation of resources by other products may result in adversarial supply. In this work, we also consider adversarial inventories. However, our paper focuses on resolving the censoring effect on demand functions and reducing regret through online learning methodologies. Perakis and Roels, (2010) also adopts regret as a metric and considers censoring effect on the demand data as we do. However, their definition of minimax and maximin regret are different from ours. Also, instead of providing theoretical solutions, the censoring effect is only considered in empirical validation to test the robustness of the control system. Other representative works on NRM including Gallego and Van Ryzin, (1994) that proposes a classic model for dynamic pricing with stochastic demand, highlighting structural monotonicity and showing that simple policies can be asymptotically optimal; Talluri and Van Ryzin, (1998), which shows bid-price controls are near-optimal in large capacities but are not strictly optimal; and Meissner and Strauss, (2012) that addresses overbooking with product-specific no-shows via a randomized linear program.
Pricing with inventory concerns
Dynamic pricing problems began to incorporate inventory constraints with the work of Besbes and Zeevi, (2009), which assumed a fixed initial stock available at the start of the selling period. They introduced near-optimal algorithms for both parametric and non-parametric demand distributions, operating under the assumption that the inventory was non-replenishable and non-perishable. Wang et al., (2014) adopted a comparable framework but allowed for customer arrivals to follow a Poisson process. In these earlier works, the actual demand is fully disclosed until the inventory is depleted. Subsequent research allowed inventory replenishment, with the seller’s decisions encompassing both pricing and restocking at each time interval. Chen et al., (2019) proposed a demand model subject to additive/multiplicative noise and developed a policy that achieved $O(\sqrt{T})$ regret. More recent studies, such as those by Chen et al., (2020); Keskin et al., (2022); Xu et al., (2025) explored the dynamic pricing of perishable goods where unsold inventory would expire. However, the uncensored demand is observable as assumed in both works. Specifically, Chen et al., (2020) allowed recouping backlogged demand, albeit at a cost, and introduced an algorithm with optimal regret. Keskin et al., (2022) focused on cases where both fulfilled demands and lost sales were observable.
Chen et al., (2021) and their subsequent work, Chen et al., (2024), are the closest works to ours as they adopt similar problem settings: In their works, the demand is censored by the inventory level and any leftover inventory or lost sales disappear at the end of each period. With the assumption of concave reward functions and the restriction of at most $m$ price changes, Chen et al., (2021) proposed MLE-based algorithms that attain a regret of $\tilde{O}(T^{\frac{1}{m+1}})$ in the well-separated case and $\tilde{O}(T^{\frac{1}{2}+\epsilon})$ for some $\epsilon=o(1)$ as $T\rightarrow\infty$ in the general case. Under similar assumptions (except infinite-order smoothness), Chen et al., (2024) developed a reward-difference estimator, with which they not only enhanced the prior result for concave reward functions to $\tilde{O}(\sqrt{T})$ but also obtained a general $\tilde{O}(T^{2/3})$ regret for non-concave reward functions. Our problem model mirrors their difficulty in lacking access to both the uncensored demand and its gradient. However, they allowed the sellers to determine inventory levels with sufficient flexibility, hence better balancing the information revealed by the censored demand and the reward from (price, inventory) decisions. On the other hand, we assume the inventory level at each time period is provided adversarially by nature, which could impede us from learning the optimal price in the worst-case scenarios. Furthermore, due to the non-stationarity of inventory levels in our setting, the optimal price $p_{t}^{*}$ deviates over time. Given this, the searching-based methods adopted in Chen et al., (2024) is no longer applicable to our problem.
3 Problem Setup
In this work, we study the following online non-contextual dynamic pricing problem. At each time step $t=1,2,\ldots,T$, the seller (we) first receives an inventory quantity $\gamma_{t}$ determined by adversarial nature. Then we propose a price $p_{t}$ for these items and receive an inventory-censored demand $D_{t}:=\min\{d_{t},\gamma_{t}\}$. Here $d_{t}=a-b\cdot p_{t}+N_{t}$ is a potential linear demand, where $a,b$ are unknown parameters and $N_{t}$ is the demand noise.
In the following subsections, we will adopt definitions and assumptions to further clarify the scope of our methodologies.
3.1 Definitions
Here we define some key quantities that are involved in the algorithm design and analysis. Firstly, there are two different types of demand functions.
Definition 3.1 (Demand functions).
Denote $d_{t}(p):=a-bp+N_{t}$ as the potential demand function, and $d(p):=a-bp$ as the expected potential demand function. Denote $D_{t}(p):=\min\{\gamma_{t},d_{t}(p)\}$ as the censored demand function.
Then we define distributional functions of the noise $N_{t}$.
Definition 3.2 (Distributional functions).
For $N_{t}$ as the demand noise, denote $F(x)$ as its cumulative distribution function (CDF), $x\in\mathbb{R}$. Also, denote
[…]
as the integrated CDF.
We will make more assumptions on the noise distribution later. Notice that we do not assume the existence of PDF for $N_{t}$. However, if there exists its PDF in specific cases, we will adopt $f(x)$ as a notation.
Finally, we may define the revenue function and the regret.
Definition 3.3 (Revenue function).
Denote $r_{t}(p)$ as the expected revenue function of price $p$, satisfying
$r_{t}(p):=p\cdot\mathbb{E}_{N_{t}}[D_{t}(p)|\gamma_{t}],p\geq 0.$ (2)
Also, denote $p_{t}^{*}:=\mathop{\mathrm{argmax}}_{p}r_{t}(p)$ as the optimal price at time $t$.
Definition 3.4 (Regret).
Denote
$Regret:=\sum_{t=1}^{T}r_{t}(p_{t}^{*})-r_{t}(p_{t})$ (3)
as the cumulative regret (or regret) of the price sequence $\{p_{t}\}_{t=1}^{T}$.
3.2 Assumptions
We make reasonable and mild assumptions as follows. Firstly, we assume boundaries for parameters and price.
Assumption 3.5 (Boundedness).
There exist known finite constants $a_{\max},b_{\min},b_{\max},\gamma_{\min},c>0$ such that $0<a\leq a_{\max}$, $0<b_{\min}\leq b\leq b_{\max}$, $\gamma_{t}\geq\gamma_{\min}$, $N_{t}\in[-c,c]$. Also, we restrict the proposed price $p_{t}$ at any $t=1,2,\ldots,T$ satisfies $0\leq p_{t}\leq p_{\max}$ with a known finite constant $p_{\max}>0$.
Secondly, we make assumptions on the noise distribution.
Assumption 3.6 (Noise Distribution).
Each $N_{t}$ is drawn from an unknown independent and identical distribution (i.i.d.) satisfying $\mathbb{E}[N_{t}]=0$. The CDF $F(x)$ is $L_{F}$-Lipschitz. Also, according to Assumption 3.5, we have $F(-c)=0,F(c)=1$.
Thirdly, we make assumptions on the inequality relationships among parameters:
Assumption 3.7 (Inequalities of Parameters).
The parameters and constants satisfy the following inequalities:
  1. 1.
    $a-c>\gamma_{t},\forall t\in[T]$. Demands at $p_{t}=0$ must be censored.
  2. 2.
    $\gamma_{t}>2c,\forall t\in[T]$. Inventory level exceeds noise support.
  3. 3.
    $a-bp_{\max}-c>0$ and therefore $p_{\max}\leq\frac{a}{b}$. Demands must be positive.
  4. 4.
    $\gamma_{\min}>a_{\max}-b_{\min}p_{\max}+c$. Demands at $p_{t}=p_{\max}$ must be uncensored. Meanwhile, we denote $\gamma_{0}:=a_{\max}-b_{\min}p_{\max}+c$ for further use.
  5. 5.
    Without loss of generality, let $p_{\max}\geq\frac{a}{2b}+1$. Optimal price must be included in $[0,p_{\max}]$.
Finally, we assume that the time horizon $T$ is sufficiently large, such that it will not confound with any constant or coefficient.
Assumption 3.8 (Large $T$).
For any given polynomial of parameters $poly(a_{\max},b_{\max},1/b_{\min},c,1/\gamma_{\min},p_{\max})$, we have $T>poly$.
4 Algorithm
In this section, we present our core algorithm,C20CB, which stands for a Closest-To-Zero Confidence Bound strategy that proposes asymptotically optimal prices over differentiated inventory levels and censoring effects. The structure of C20CB is displayed as Algorithm 1.
Algorithm 1 C20CB:Closest-To-Zero Confidence Bound
1:  Input: Quantities $\tau$ and $a_{\max},b_{\min},b_{\max},p_{\max},c$, coefficients $\gamma_{0},C_{a},C_{b},C_{F},C_{G},C_{N},C_{\tau}$.
2:  STAGE 1: Pure Exploration
3:  for  $t=1,2,\ldots,\tau$ do
4:     Sample and propose a price $p_{t}\sim U[0,p_{\max}]$ uniformly at random.
5:     Observe demand $D_{t}$ and indicators $e_{i,t}:=\mathds{1}[D_{t}\geq\frac{i\cdot\gamma_{t}+(4-i)\gamma_{0}}{4}],i=1,2,3$.
6:  end for
7:  Estimate
$\hat{b}=$ $\frac{1}{4p_{\max}\cdot\frac{1}{\tau}\sum_{t=1}^{\tau}\frac{e_{1,t}-e_{2,t}}{\gamma_{t}-\gamma_{0}}}$ (4)
$\hat{a}=$ $\frac{1}{\tau}\sum_{t=1}^{\tau}(\hat{b}p_{\max}e_{3,t}+\frac{3\gamma_{t}+\gamma_{0}}{4}).$
8:  STAGE 2: Optimistic Acting
9:  Denote $\Delta:=(C_{a}+C_{b}\cdot p_{\max})\cdot\frac{1}{\sqrt{\tau}}$.
10:  Define $M:=\lfloor\frac{c}{2\Delta}\rfloor$ and $w_{k}:=2k\Delta,k=-M,-M+1,\ldots,-1,0,1,\ldots,M$.
11:  for $t=1,2,\ldots,2M+1$ do
12:     Let $k_{t}:=-M-1+t$ and propose $p_{t}=\frac{w_{k_{t}}-(\gamma_{t}-\hat{a})}{\hat{b}}$.
13:     Observe $D_{t}$ and $\mathds{1}_{t}:=\mathds{1}[D_{t}<\gamma_{t}]$.
14:     Initialize $F_{k_{t}}\leftarrow\mathds{1}_{t},N_{k_{t}}\leftarrow 1,G_{k_{t}}\leftarrow D_{t}-\gamma_{t}+c,\Delta_{k_{t}}\leftarrow C_{F}\cdot b_{\max}p_{\max}+C_{G}+C_{b}\cdot\frac{1}{\sqrt{\tau}}$.
15:  end for
16:  for $t=1,2,\ldots,T-\tau-(2M+1)$ do
17:     if $\gamma_{t}\geq\frac{\hat{a}+C_{a}\cdot\frac{1}{\sqrt{\tau}}}{2}+c$ then
18:        Propose $p_{t}=\frac{\hat{a}}{2\hat{b}}$ and continue to $t+1$ (without recording feedback).
19:     end if
20:     Initialize $k_{t}\leftarrow M,\rho_{t}\leftarrow+\infty$.
21:     for $k=M,M-1,\ldots,-M+1,-M$ do
22:        Denote $p_{k,t}:=\frac{w_{k}-(\gamma_{t}-\hat{a})}{\hat{b}}$ and $\hat{r}_{k,t}:=\gamma_{0}-c+G_{k}-\hat{b}\cdot p_{k,t}\cdot F_{k}$.
23:        if $\hat{r}_{k,t}-\Delta_{k}\leq 0\leq\hat{r}_{k,t}+\Delta_{k}$ then
24:           Update $k_{t}\leftarrow k,\rho_{t}\leftarrow 0$, and Break.
25:        end if
26:        Let $\rho_{k,t}:=\min\{|\hat{r}_{k,t}-\Delta_{k}|,|\hat{r}_{k,t}+\Delta_{k}|\}$.
27:        if $\rho_{k,t}<\rho_{t}$ then
28:           Update $\rho_{t}\leftarrow\rho_{k,t}$ and $k_{t}\leftarrow k$.
29:        end if
30:     end for
31:     if $\hat{r}_{k,t}-\Delta_{k}>0,\forall k=-M,-M+1,\ldots,M-1,M$ then
32:        Propose $p_{t}=\frac{\hat{a}}{2\hat{b}}.$
33:     else
34:        Propose $p_{t}=p_{k_{t},t}$.
35:     end if
36:     Observe $D_{t}$ and $\mathds{1}_{t}=\mathds{1}[D_{t}<\gamma_{t}]$.
37:     Update
$F_{k_{t}}\leftarrow\frac{N_{k_{t}}F_{k_{t}}+\mathds{1}_{t}}{N_{k_{t}}+1},\ G_{k_{t}}\leftarrow\frac{N_{k_{t}}G_{k_{t}}+D_{t}-\gamma_{t}+c}{N_{k_{t}}+1},\ \Delta_{k_{t}}\leftarrow\frac{C_{N}}{\sqrt{N_{k_{t}}+1}}+\frac{C_{\tau}}{\sqrt{\tau}}$ (5)
38:     Update $N_{k_{t}}\leftarrow N_{k_{t}}+1$.
39:  end for
4.1 Algorithm Design Overview
Our algorithm has two stages:
  1. 1.
    Stage 1: Exploration: During the first $\tau=\sqrt{T}$ rounds, the seller (we) proposes uniformly random prices in the range of $[0,p_{\max}]$. By the end of Stage 1, we obtain $\hat{a}$ and $\hat{b}$ as plug-in estimators of $a$ and $b$ in the following stage.
  2. 2.
    Stage 2: Optimistic Decision: We estimate the derivatives of the revenue function at discretized prices $\{p_{k,t}\}$’s. For each $p_{k,t}$, we not only estimate $r^{\prime}_{t}(p_{k,t}$ but also maintain an error bar of that estimate. At each time $t$, we propose the price whose corresponding error bar covers 00 or closest to 00 if no covering exists.
Algorithm 1 exhibits several advantageous properties. It is suitable for processing streaming data as the constructions of $\hat{a},\hat{b},\hat{r}^{\prime}_{t}(\cdot)$ are updated incrementally with each new observation (including $e_{i,t},D_{t},\mathds{1}_{t}$) without the need of revisiting any historical data. Additionally, it consumes $\tilde{O}(T^{\frac{5}{4}})$ time complexity and $O(T^{\frac{1}{4}})$ extra space, which are plausible for large $T$. A potential risk of computation might arise on the calculation of $\hat{b}$, where $\sum_{t=1}^{\tau}e_{1,t}-e_{2,t}$ can be 0 with a small but nonzero probability. Although this event does not undermine the high-probability regret guarantee, it might still be harmful to the computational system for numerical experiments. To mitigate this incident in practice, we may either extend Stage 1 until one non-zero $e_{1,t}-e_{2,t}=1$ is observed, or restart Stage 1 at $t=\tau$.
In the following sections, we will introduce each stage and technical component in details.
4.2 Pure-Exploration to Estimate Parameters from Biased Observations
We incorporate a uniform-exploration phase for estimating $a$ and $b$ in our algorithm, bypassing the obstacle brought by demand censoring. This approach is supported by the following insight: When $Y$ is a uniformly distributed random variable within a closed interval $[L,R]$, and $X$ is another random variable, independent to $Y$ and also distributed within $[L,R]$, we have:
$\mathbb{E}[\mathds{1}[Y\geq X]]=\Pr[Y\geq X]=\mathbb{E}[\Pr[Y\geq X|X]]=\mathbb{E}[\frac{X-L}{R-L}]=\frac{\mathbb{E}[X]-L}{R-L}.$ (6)
Here the second step uses the Law of Total Expectation. Eq. 6 indicates that we can derive an unbiased estimator of $\mathbb{E}[X]$ through $\mathds{1}[Y\geq X]$ even in the absence of any direct observation of $X$. Looking back to our algorithm, when $p_{t}\sim U[0,p_{\max}]$, we have
$\mathbb{E}[e_{i,t}]$ $=\mathbb{E}[\mathds{1}[a-bp_{t}+N_{t}\geq\gamma_{i}]]=\mathbb{E}[\mathbb{E}[N_{t}\geq\gamma_{i}-a+bp_{t}|N_{t}]]=\mathbb{E}[\frac{N_{t}-\gamma_{i}+a}{bp_{\max}}]=\frac{a-\gamma_{i}}{bp_{\max}}.$ (7)
Here $\gamma_{i}:=\frac{i\gamma_{t}+(4-i)\gamma_{0}}{4}$. The last equality comes from $\mathbb{E}[N_{t}]=0$. By deploying different $\gamma_{i}$ at $i=1,2,3$, we can estimate $a$ and $b$ through the observations of $e_{i,t}$, effectively circumventing the censoring effect. A similar technique has been utilized by Fan et al., (2024) to construct an unbiased estimator of the valuations instead of the demands as we concern. However, their application of uniform exploration might be sub-optimal as they adopt an exploration-then-exploitation design in each epoch. On the contrary, our algorithm uses this uniform exploration merely as a trigger of further learning. Our tight regret bound indicates that uniform exploration can still contribute to an optimal algorithm for a broad range of online learning instances.
4.3 Optimistic Strategy to Balance Derivatives Estimates v.s. Revenue Loss
Figure 1: The price which C20CB proposes based on confidence bounds of $\hat{r}_{k,t}$: (a) If there exist prices whose error bar contain 00, then we propose the largest price among them. (b) If no error bar contains 00 but there does exist at least one below 00, we propose the price whose corresponding error bar is closest to 00. (c) If all error bars are above 00, then we propose $\frac{\hat{a}}{2\hat{b}}$ for pure exploitation.
We proceed to Stage 2 with $\hat{a}$ and $\hat{b}$ established, which indicates we have an estimate on the expected potential linear demand $d_{t}(p)=a-bp$. However, we are still unaware of the noise distribution, which is crucial for the current optimal price as the inventory level $\gamma_{t}$ partially censors the noise.
In order to balance the learning of noise distribution versus the loss of proposing sub-optimal prices, we adopt an optimistic which involves the following components:
  1. (i)
    We discretize the $[-c,c]$ domain of noise CDF $F(\cdot)$ and its intergration $G(\cdot)$ into small intervals of length $2\Delta$, with $\Delta=O(\frac{1}{T^{1/4}})$. At the center of each interval $k$ (which is $2k\Delta$) we maintain independent estimates of $F$ and $G$, including their expectations and high-confidence error bars.
  2. (ii)
    At each time $t$, we construct a set of discrete prices $\{p_{k,t}\}_{k=-M}^{M}$ such that the quantity $\gamma_{t}-\hat{a}+\hat{b}p_{k,t}$ matches the center of the interval $k$ established above. According to Eq. 13, we construct an estimate of each $r^{\prime}_{t}(p_{k,t})$ with plug-in estimators $\hat{a},\hat{b}$ and discrete estimators of $F$ and $G$ for the specific Interval $k$. Again, the estimate includes its expectation and error bar.
  3. (iii)
    Since the optimal price $p_{t}^{*}$ satisfies $r^{\prime}_{t}(p_{t}^{*})=0$, we identify the discrete price $p_{k,t}$ where the derivative estimate is "possibly 00" or "closest to 00". We illustrate this process in Figure 1, which includes the following three cases:
    1. (a)
      If there exists some $p_{k,t}$ such that the corresponding error bar (of its derivative estimate) contains 00, we propose the largest price satisfying this condition.
    2. (b)
      If there does not exist $p_{k,t}$ whose corresponding error bar contains 00 (but there does exist error bar below 00), we propose the price whose error bound is closest to 00.
    3. (c)
      If all error bars are above 00, indicating that the reward function is monotonically increasing over the “censoring area”, we propose $p_{t}=\frac{\hat{a}}{2\hat{b}}$ to exploit the non-censoring optimal price $\frac{a}{2b}$. In this case, we do not need to record any observations nor to update any parameter/estimate.
  4. (iv)
    After proposing the price $p_{k_{t},t}$ and observing feedback $D_{t}$ and $\mathds{1}_{t}$, we update the estimates of $F(\cdot)$ and $G(\cdot)$ for Interval $k_{t}$ in which $\gamma_{t}-\hat{a}+\hat{b}p_{k_{t},t}$ exists.
In a nutshell, we maintain estimates and error bars of $F(\cdot)$ and $G(\cdot)$ at discrete points $2k\Delta$, and map each $2k\Delta$ to a corresponding price $p_{k,t}$ once an inventory $\gamma_{t}$ occurs. Then we propose the price whose derivative estimate $\hat{r}_{k,t}\pm\Delta_{k}$ is closest-to-zero among all $k$. Finally, we update the estimates with observations.
Here we provide an intuition of the optimality: On the one hand, the width of interval can tolerate the error of mapping from $2k\Delta$ to $p_{k,t}$, and the Lipschitzness of $F$ and $G$ ensures that our estimate within each small interval is roughly correct. On the other hand, we can show that the closest-to-zero derivative estimate implies a closest-to-$p_{t}^{*}$ price according to some locally strong convexity in the neighborhood of $p_{t}^{*}$. As we have smoothness on the regret function, we suffer a quadratic loss at $(T\cdot\frac{1}{T^{1/4}})^{2}=O(\sqrt{T})$ cumulatively, which balances the loss of Stage 1 that costs $O(\tau)=O(\sqrt{T})$ as well. For a rigorous regret analysis, we kindly refer the readers to Section 5.
4.4 Technical Novelty.
To the best of our knowledge, we are the first to introduce optimism on the derivatives and achieve optimal regret in an adversarial online learning problem. In contrast, existing works either develop optimistic algorithms on the reward (or loss) function as the original UCB strategy (Lai and Robbins,, 1985), or instead use unbiased stochastic gradients and conduct first-order methods for online optimization (Hazan,, 2016).
5 Regret Analysis
In this section, we analyze the cumulative regret of our algorithm and show a $\tilde{O}(\sqrt{T})$ regret guarantee with high probability. We leave all of the proof details to Appendix A, and here we only display key lemmas and proof sketches.
We firstly propose our main theorem.
Theorem 5.1 (Regret).
Let $\tau=\frac{1}{\sqrt{T}}$ in Algorithm 1. For any adversarial $\{\gamma_{t}\}_{t=1}^{T}$ input sequence, C20CB suffers at most $\tilde{O}\left(\sqrt{T}\cdot\log\frac{T}{\delta}\right)$ regret, with probability $\Pr\geq 1-\delta$.
Proof.
In order to prove Theorem 5.1, we have to show the following three components:
  1. 1.
    The reward function $r_{t}(p)$ is unimodal. Also, $r_{t}(p)$ is smooth at $p_{t}^{*}$, and is strongly concave on a neighborhood of $p_{t}^{*}$.
  2. 2.
    The estimation error of $a$ and $b$ are bounded by $O(\frac{1}{T^{1/4}})$ at the end of Stage 1.
  3. 3.
    The price whose derivative estimate has the closest-to-zero confidence bound is asymptotically close to $p_{t}^{*}$.
In the following, we present each corresponding lemma regarding to the roadmap listed above.
Lemma 5.2 (revenue function $r_{t}(p)$).
For the expected revenue function $r_{t}(p)$ defined in eq. 2, the following properties hold:
  1. 1.
    We have
    $r_{t}(p)$ $=p(\gamma_{t}-c+G(c)-G(\gamma_{t}-a+bp))$ (8)
    $r^{\prime}_{t}(p)$ $=\gamma_{t}-c+G(c)-G(\gamma_{t}-a+bp)-bp\cdot F(\gamma_{t}-a+bp).$
  2. 2.
    There exists a constant $L_{r}>0$ such that $r^{\prime}(p)$ is $L_{r}$-Lipschitz.
  3. 3.
    $r^{\prime}(p)$ is monotonically non-increasing.
  4. 4.
    $r_{t}(p)$ is unimodal: There exists a unique $p_{t}^{*}\in[0,\frac{a}{b}]$ such that $r^{\prime}_{t}(p_{t}^{*})=0$, and $r_{t}(p)$ monotonically increase in $[0,p_{t}^{*}]$ and decrease in $[p_{t}^{*},\frac{a}{b}]$. Notice that $\frac{a}{b}>p_{\max}$ according to Assumption 3.7.
  5. 5.
    $r_{t}(p)$ is smooth at $p_{t}^{*}$: There exists a constant $C_{s}>0$ such that $r_{t}(p_{t}^{*})-r_{t}(p)\leq C_{s}(p_{t}^{*}-p)^{2},\forall p\in[0,p_{\max}]$.
  6. 6.
    $r_{t}(p)$ is locally strongly concave: There exist $\epsilon_{t}>0$ and $C_{\epsilon}>0$ such that $\forall p_{1},p_{2}\in[p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t}]$ we have $|r^{\prime}_{t}(p_{1})-r^{\prime}_{t}(p_{2})|\geq C_{\epsilon}\cdot|p_{1}-p_{2}|$.
  7. 7.
    There exists a constant $C_{v}>0$ such that for any $t\in[T]$ and $p\in(p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t})$, we have $|r_{t}(p_{t}^{*})-r_{t}(p)|\leq C_{v}\cdot(r^{\prime}_{t}(p))^{2}$.
Please kindly check Section A.1 as a detailed proof of Lemma 5.2. The properties of $r_{t}(p)$ and $r^{\prime}_{t}(p)$ enable us upper bound the cost of estimation error and decision bias. In the following, we propose a lemma that serves as a milestone of estimation error upper bounds.
Lemma 5.3 (Estimation error of $a$ and $b$).
With probability $\Pr\geq 1-2\eta\delta$, we have
$|\hat{b}-b|\leq$ $C_{b}\cdot\frac{1}{\sqrt{\tau}},$ (9)
$|\hat{a}-a|\leq$ $C_{a}\cdot\frac{1}{\sqrt{\tau}},$
where $C_{a}:=p_{\max}(C_{b}+b_{\max}\cdot\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}})$ and $C_{b}:=\frac{8b_{\max}^{2}}{\gamma_{\min}-\gamma_{0}}\cdot\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}$.
The key observation to prove this lemma lies in the expectation of each $e_{i,t}$ that indicates whether the demand exceeds certain level under uniformly distributed prices. With the help of a method-of-moment estimate as introduced in Section 4.2, we get rid of the influence of noise distribution and achieves an unbiased estimator of $\frac{1}{b}$ (and therefore $\hat{b}$). With $\hat{b}$ serving as a plug-in estimator, we later get $\hat{a}$. By applying Hoeffding’s Inequalities, we obtain those error bounds. We defer the detailed proof of Lemma 5.3 to Section A.2. With the help of Lemma 5.3, we may upper bound the estimation error of $r^{\prime}_{t}(p)$ at discrete prices. The error bound is displayed as the following lemma.
Lemma 5.4 (Estimation error of $r^{\prime}_{t}(p_{k,t})$).
There exists constants $C_{N}>0,C_{\tau}>0$ such that for any $t\in[T],k\in\{-M,-M+1,\ldots,M-1,M\}$, its holds with probability $\Pr\geq 1-6\eta\delta$
$|r^{\prime}_{t}(p_{k,t})-\hat{r}_{k,t}|\leq C_{N}\cdot\frac{1}{\sqrt{N_{k}(t)}}+C_{\tau}\cdot\frac{1}{\sqrt{\tau}}=\Delta_{k}(t).$ (10)
Here $N_{k}(t)$ and $\Delta_{k}(t)$ denotes the value of $N_{k}\text{ and }\Delta_{k}$ at the beginning of time period $t$.
We kindly refer the readers to Section A.3 as a rigorous proof of Lemma 5.4. Given this lemma, the derivatives of each discrete price $p_{k,t}$ is truthfully reflected by their corresponding error bound. Therefore, we intuitively see that the closest-to-zero confidence bound represents the closest-to-$p_{t}^{*}$ discrete price. We formulate this intuition as the following lemma:
Lemma 5.5 (Closest-to-zero confidence to performance).
Denote $\Delta_{k}(t)$ as the value of $\Delta_{k}$ at the beginning of period $t$. There exists two constants $N_{0}>0,N_{1}>0$ such that for any $t=1,2,\ldots,T$ in Stage 2, either of the following events occurs with high probability.
  1. 1.
    When $\exists k\in\{-M,-M+1,\ldots,M-1,M\}$ such that the Number $k$ confidence bound satisfies $\hat{r}_{k,t}-\Delta_{k}(t)\leq 0\leq\hat{r}_{k,t}+\Delta_{k}(t)]$, and also $N_{k}(t)>N_{0}$, then we have $p_{k,t}\in[p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t}]$. Furthermore, there exists constant $C_{in}$ such that $r_{t}(p_{t}^{*})-r_{t}(p_{k,t})\leq C_{in}(\frac{1}{N_{k}(t)}+\frac{1}{\tau})$.
  2. 2.
    When there exists no confidence bound that contains 00, i.e. either $\hat{r}_{k,t}-\Delta_{k}(t)>0$ or $\hat{r}_{k,t}+\Delta_{k}(t)<0,\forall k\in\{-M,-M+1,\ldots,M-1,M\}$ (happens at least for one $k$), and also $N_{k}(t)>N_{1}$, then we have
    $\inf_{k}\min\{|\hat{r}_{k,t}-\Delta_{k}(t)|,|\hat{r}_{k,t}+\Delta_{k}(t)|\}\leq\frac{L_{r}(C_{a}+C_{b}\cdot p_{\max})}{2b_{\min}}\cdot\frac{1}{\sqrt{\tau}},$
    and also $p_{k,t}\in[p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t}]$ Furthermore, there exists constant $C_{out}$ such that $r_{t}(p_{t}^{*})-r_{t}(p_{k,t})\leq C_{out}(\frac{1}{N_{k}(t)}+\frac{1}{\tau})$.
The intuition to prove Lemma 5.5 is twofold:
  1. 1.
    When an error bar contains 00, the true derivative of the corresponding price is close to 00 within the distance of its error bound. By applying Lemma 5.2 Property (7), we may upper bound the performance loss with the square of its derivatives, which is further upper bounded by the square of error bound.
  2. 2.
    When no error bar contains 00, there exists an adjacent pair of prices whose error bars are separated by $y=0$. On the one hand, their derivatives difference is upper bounded due to the Lipschitzness of $r^{\prime}_{t}(p)$. On the other hand, the same derivatives difference is lower bounded by the closest-to-zero confidence bound. Therefore, the gap between $y=0$ and the closest-to-zero confidence bound should be very small, and we still have a comparably small $|r^{\prime}_{t}(p_{t})|$ if $p_{t}$ possesses that confidence bound. As a consequence, we have similar upper bound on the performance loss comparing with Case (1), up to constant coefficients.
The detailed proof of Lemma 5.5 is presented in Section A.4. Finally, we have a lemma that upper bounds the regret of proposing $p_{t}=\frac{\hat{a}}{2\hat{b}}$ under specific circumstances.
Lemma 5.6 (Proposing $\frac{\hat{a}}{2\hat{b}}$).
When $\gamma_{t}>\frac{\hat{a}+C_{a}\cdot\frac{1}{\sqrt{\tau}}}{2}+c$ and when $\hat{r}_{k,t}-\Delta_{k}(t)>0,\forall k=-M,-M+1,\ldots,M-1,M$, we have $p_{t}^{*}=\frac{a}{2b}$ and there exists a constant $C_{non}$ such that
$r_{t}(\frac{a}{2b})-r_{t}(\frac{\hat{a}}{\hat{2}b})\leq C_{non}\frac{1}{\tau}.$ (11)
The intuition of Lemma 5.6 is that $\frac{a}{2b}$ is the optimal price without censoring, and we only need to show that either the optimal price or $\frac{a}{2b}$ is not censored (which are equivalent as the revenue function is unimodal). We defer its proof to Section A.5. This lemma serves as the last puzzle of the proof. With all lemmas above, we may upper bound the cumulative regret as follows:
$Regret$ $=\sum_{t=1}^{T}r_{t}(p_{t}^{*})-r_{t}(p)$ (12)
$\leq\tau\cdot a_{\max}p_{\max}+(2M+1)(1+N_{0}+N_{1})\cdot a_{\max}p_{\max}$
$\quad+\sum_{t=1}^{T}(\max\{C_{in},C_{out}\})(\frac{1}{N_{k}(t)}+\frac{1}{\tau})+\sum_{t=1}^{T}C_{non}\cdot\frac{1}{\tau}$
[…]
$=\tilde{O}(\sqrt{T}+\sum_{k=1}^{2M+1}\sum_{i_{k}=1}^{N_{k}(T)}\frac{1}{i_{k}})$
$=\tilde{O}(\sqrt{T}+T^{\frac{1}{4}}\log T)$
$=\tilde{O}(\sqrt{T}).$
By applying a union bound, we know that Eq. 12 holds with probability
$\Pr\geq 1-2\eta\delta-6\eta\delta\cdot T(2M+1)\geq 1-20\frac{c}{2(C_{a}+C_{b}\cdot p_{\max})}T^{5/4}\cdot\eta\delta.$
Here the first part comes from Lemma 5.3, and the second part comes from Lemma 5.4 for any $t\in[T]$ and $k\in\{-M,\ldots,M\}$. Let $\eta:=\frac{C_{a}+C_{b}\cdot p_{\max}}{10c\cdot T^{5/4}}$, and we show that Theorem 5.1 holds. ∎
Remark 5.7.
This $\tilde{O}(\sqrt{T})$ regret upper bound is (near) optimal up to $\log{T}$ factors, as it matches the $\Omega(\sqrt{T})$ information-theoretic lower bound proposed by Broder and Rusmevichientong, (2012) for a no-censoring problem setting with linear noisy demand.
From the analysis above, we notice that the threshold of learning the optimal prices in our problem setting is still the estimation of parameters. The assumptions we made in Section 3 scales the efficiency of learning $\hat{b}$ and $\hat{a}$, which we will discuss in Section 6.
6 Discussions
Here we discuss the limitations, potential extensions and impacts of our work.
Generalization to Unbounded Noises
We assume the noise is bounded in a constant-width range. This assumption streamlines the pure-exploration phase and facilitates the estimation of the parameters $b$ and $a$. While our methods and results can be extended to unbounded $O(\frac{1}{\log T})$-subGaussian noises by simple truncation, challenges remain for handling generic unbounded noises. Moreover, the problem can be more sophisticated with dual-censoring, both from above by inventory–as we have discussed– and from below by 00, especially when considering unbounded noises.
Extensions to Contextual Pricing
In this work, we assume $a$ and $b$ are static, which may not hold in many real scenarios. Example 1.1 serves as a good instance, showcasing significant fluctuations in popularity across different performances. A reasonable extension of our work would be modeling $a$ and $b$ as contextual parameters. Similar modelings have been adopt by Wang et al., (2025) and Ban and Keskin, (2021) in the realm of personalized pricing research.
Extensions to Non-Lipschitz Noise CDF
In this work, we assume the noise CDF as a Lipschitz function as many pricing-related works did (Fan et al.,, 2024; Tullii et al.,, 2024). This assumption enables the local smoothness at $p_{t}^{*}$ and leads to a quadratic loss. However, this prevent us from applying our algorithm to non-Lipschitz settings, which even includes the noise-free setting. In fact, although we believe that a better regret rate exists for the noise-free setting, we have to state that the hardness of problems are completely different with Lipschitz noises versus without it. Although a Lipschitz noise makes the observation “more blur”, it also makes the revenue curve “more smooth”. We would like to present an analog example from the feature-based dynamic pricing problem: When the Gaussian noise $\mathcal{N}(0,\sigma^{2})$ is either negligible (with $\sigma<\frac{1}{T}$, see Cohen et al., (2020)) or super significant (with $\sigma>1$, see Xu and Wang, (2021)), the minimax regret is $O(\log T)$. However, existing works can only achieve $O(\sqrt{T})$ regret when $\sigma\in[\frac{1}{T},1]$. We look forward to future research on our problem setting once getting rid of the Lipschitzness assumption.
Extensions to Non-linear Demand Curve
In this work, we adopt a linear-and-noisy model for the potential demands, which is standard in dynamic pricing literature. Also, we utilize the unimodal property brought by this linear demand model, even after the censoring effect is imposed. If we would like to generalize our methodologies to non-linear demand functions, especially where the unimodality does not hold any longer, we have to carefully distinguish between local optimals and saddle-points that may also cause $r^{\prime}_{t}(p)=0$ for some sub-optimal $p$. We conjecture an $\Omega(T^{\frac{m+1}{2m+1}})$ lower bound in that case, where $m$ is the time of smoothness. However, it is still unclear whether the censoring effect will introduce new local optimals or swipe off existing ones in multimodal settings.
Societal Impacts
Our research primarily addresses a non-contextual pricing model that does not incorporate personal or group-specific data, thereby adhering to conventional fairness standards relating to temporal, group, demand, and utility discrepancies as outlined by Cohen et al., (2022); Chen et al., (2023); Xu et al., (2023). However, the non-stationarity of inventory levels could result in varying fulfillment rate over time, i.e. the proportions of satisfied demands at $\{p_{t}^{*}\}$’s might be different for $t=1,2,\ldots,T$. This raises concern regarding unfairness in fulfillment rate (Spiliotopoulou and Conte,, 2022) particularly on product of significant social and individual importance.
7 Conclusions
In this paper, we studied the online pricing problem with adversarial inventory constraints imposed over time series. We introduced an optimistic strategy and a C20CB algorithm that is capable of approaching the optimal prices from inventory-censored demands. Our algorithm enjoys a regret guarantee at $\tilde{O}(\sqrt{T})$ with high probability, which is information-theoretically optimal. To the best of our knowledge, we are the first to address this adversarial-inventory pricing problem, and our results indicate that the demand-censoring effect does not substantially increase the hardness of pricing in terms of minimax regret.
Appendix A Proof Details
A.1 Proof of Lemma 5.2
Proof.
We prove each property sequentially.
  1. 1.
    For $r_{t}(p)$, we have:
    $r_{t}(p)$ $=\mathbb{E}[p_{t}\cdot D_{t}|p_{t}=p]$ (13)
    $=p\cdot\mathbb{E}[\min\{\gamma_{t},a-bp_{t}+N_{t}\}|p_{t}=p]$
    $=p\cdot\mathbb{E}[\mathds{1}[a-bp+N_{t}\leq\gamma_{t}]\cdot(a-bp+N_{t})+\mathds{1}[a-bp+N_{t}>\gamma_{t}]\cdot\gamma_{t}]$
    $=p\cdot\mathbb{E}[\mathds{1}[N_{t}\leq\gamma_{t}-a+bp]\cdot(a-bp+N_{t})+\mathds{1}[N_{t}>\gamma_{t}-a+bp]\cdot\gamma_{t}]$
[…]
    $=p(a-bp)+0+p(\gamma_{t}-a+bp)(1-F(\gamma_{t}-a+bp))-p\cdot(xF(x)-G(x))|_{\gamma_{t}-a+bp}^{c}$
    $=p\gamma_{t}-p(\gamma_{t}-a+bp)F(\gamma_{t}-a+bp)-p(c-G(c)-F(\gamma_{t}-a+bp)\cdot(\gamma_{t}-a+bp)+G(\gamma_{t}-a+bp))$
    $=p(\gamma_{t}-c+G(c)-G(\gamma_{t}-a+bp)).$
[…]
  2. 2.
    As we have assumed, $F(x)$ is $L_{F}$-Lipschitz, and therefore $F(\gamma_{t}-a+bp)$ is $b_{\max}L_{F}$-Lipschitz, and $bpF(\gamma_{t}-a+bp)$ is $(b_{\max}+b_{\max}^{2}p_{\max}L_{F})$-Lipschitz. Also, we have $\frac{dG(\gamma_{t}-a+bp)}{dp}=b\cdot F(\gamma_{t}-a+bp)\in[0,b_{\max}]$. Let $L_{r}:=(2b_{\max}+b_{\max}^{2}p_{\max}L_{F})$, and we know that $r^{\prime}_{t}(p)$ is $L_{r}$-Lipschitz.
  3. 3.
    On the one hand, we have
    $\frac{d(\gamma_{t}-c+G(c)-G(\gamma_{t}-a+bp))}{dp}=-bF(\gamma_{t}-a+bp)\leq 0.$ (14)
    On the other hand, for any $\Delta_{p}>0$, we have
    $-[b(p+\Delta_{p})\cdot F(\gamma_{t}-a+b(p+\Delta_{p}))]-[-(bpF(\gamma_{t}-a+bp))]$ (15)
    $=$ $-b\Delta_{p}F(\gamma_{t}-a+b(p+\Delta_{p}))+bp(F(\gamma_{t}-a+bp)-F(\gamma_{t}-a+b(p+\Delta_{p})))$
    $\leq$ $0+0=0$
    Since $r^{\prime}_{t}(p)=\gamma_{t}-c+G(c)-G(\gamma_{t}-a+bp)-bpF(\gamma_{t}-a+bp)$, we know that both components are monotonically non-increasing.
  4. 4.
[…]
    $r^{\prime}_{t}(0)$ $=\gamma_{t}-c+G(c)-G(\gamma_{t}-a)$ (16)
    $>\gamma_{t}-c+G(c)-G(-c)$
    $>0.$
    $r^{\prime}_{t}(\frac{a}{b})$ $=\gamma_{t}-C+G(c)-G(\gamma_{t})-b\cdot\frac{a}{b}\cdot F(\gamma_{t})$
    $=\gamma_{t}-c+G(c)-G(c+(\gamma_{t}-c))-a\cdot 1$
    $=\gamma_{t}-c+G(c)-G(c)-(G(c)+(\gamma_{t}-c))-a$
    $=\gamma_{t}-c+0-(\gamma_{t}-c)-a$
    $=-a<0.$
    Also, $r^{\prime}_{t}(p)$ is Lipschitz as we proved above. Therefore, $\exists p_{t}^{*}\in(0,\frac{a}{b})$ such that $r^{\prime}_{t}(p_{t}^{*})=0$.
    Now we show the uniqueness of $p_{t}^{*}$. If there exists $0<p_{t}^{*}<q_{t}^{*}<\frac{a}{b}$ such that $r^{\prime}_{t}(p_{t}^{*})=r^{\prime}_{t}(q_{t}^{*})=0$, then it leads to.
    $r^{\prime}_{t}(p)\equiv 0,\forall p\in[p_{t}^{*},q_{t}^{*}]\text{ due to the monotonicity of }r^{\prime}_{t}(p)$ (17)
    $\Rightarrow$ $r^{\prime}_{t}(p)\text{ is differentiable in}(p_{t}^{*},q_{t}^{*})$
    $\Rightarrow$ $r^{\prime\prime}_{t}(p)\equiv 0,\forall p\in(p_{t}^{*},q_{t}^{*})$
    $\Rightarrow$ $F(\gamma_{t}-a+bp)\text{ is differentiable with respect to}p\text{ on }(p_{t}^{*},q_{t}^{*})$
    $\Rightarrow$ $f(\gamma_{t}-a+bp)\text{ exists on }(p_{t}^{*},q_{t}^{*})$
    $\Rightarrow$ $-2bF(\gamma_{t}-a+bp)-b^{2}pf(\gamma_{t}-a+bp)\equiv 0,\forall p\in(p_{t}^{*},q_{t}^{*})$
[…]
    $\Rightarrow$ $0\displaystyle 0\leq F(\gamma_{t}-a+bp_{t}^{*})\leq\lim_{p\rightarrow q_{t}^{*}-}F(\gamma_{t}-a+bp)=$ 0
    $\Rightarrow$ $F(\omega)\equiv 0,\forall\omega<\gamma_{t}-a+bq_{t}^{*}$
    $\Rightarrow$ $r^{\prime}_{t}(p_{t}^{*})=\gamma_{t}-c+G(c)-G(\gamma_{t}-a+bp_{t}^{*})-bp_{t}^{*}\cdot F(\gamma_{t}-a+bp_{t}^{*})$
[…]
    $\qquad\quad=\gamma_{t}-c+G(c)$
    $\qquad\quad\geq\gamma_{t}-c$
    $\qquad\quad>0.$
    This leads to contradictions that $r^{\prime}_{t}(p_{t}^{*})=0$. Therefore, $p_{t}^{*}$ is unique. Given this, we know that $r_{t}(p)$ is unimodal, which increases on $(0,p_{t}^{*})$ and decreases on $(p_{t}^{*},\frac{a}{b})$.
  5. 5.
    Since $p_{t}^{*}$ is unique, and $r^{\prime}_{t}(p)$ is $L_{r}$-Lipschitz, we have:
    $r_{t}(p_{t}^{*})-r_{t}(p)\leq\frac{L_{r}}{2}\cdot(p_{t}^{*}-p)^{2},\forall p\in[0,\frac{a}{b}].$ (18)
  6. 6.
    From the proof of part (4), we know that $F(\gamma_{t}-a+bp_{t}^{*})>0$ (or otherwise $r^{\prime}_{t}(p_{t}^{*})>0$ leading to contradiction). Denote $\epsilon_{t}:=\frac{F(\gamma_{t}-a+bp_{t}^{*})}{2L_{F}b_{\max}}$, and we have:
    $F(\gamma_{t}-a+b(p_{t}^{*}-\epsilon_{t}))\geq$ $F(\gamma_{t}-a+bp_{t}^{*})-L_{F}\cdot b\cdot\epsilon_{t}$ (19)
    $\geq$ $F(\gamma_{t}-a+bp_{t}^{*})-L_{F}\cdot b\cdot\frac{F(\gamma_{t}-a+bp_{t}^{*})}{2L_{F}\cdot b_{\max}}$
    $=$ $\frac{F(\gamma_{t}-a+bp_{t}^{*})}{2}.$
    Let $C_{\epsilon}:=\frac{b_{\min}}{2}\cdot\inf_{\gamma_{t}\in[\gamma_{\min},\gamma_{\max}]}F(\gamma_{t}-a+bp_{t}^{*})$. As $[\gamma_{\min},\gamma_{\max}]$ is a close set and $F(\gamma_{t}-a+bp_{t}^{*})$ holds for any $\gamma_{t}\in[\gamma_{\min},\gamma_{\max}]$, we know that $C_{\epsilon}>0$ is a universal constant. Given this coefficient, for any $p_{1},p_{2}\in[p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t}],p_{1}<p_{2}$, we have
    $r^{\prime}_{t}(p_{1})-r^{\prime}_{t}(p_{2})$ (20)
    $=$ $-(G(\gamma_{t}-a+bp_{1})-G(\gamma_{t}-a+bp_{2}))-(bp_{1}F(\gamma_{t}-a+bp_{1})-bp_{2}F(\gamma_{t}-a+bp_{2}))$
    $\geq$ $-(G(\gamma_{t}-a+bp_{1})-G(\gamma_{t}-a+bp_{2}))$
[…]
    $\geq$ $\min_{\omega\in[p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t}]}F(\omega)\cdot b(p_{2}-p_{1})$
    $\geq$ $F(\gamma_{t}-a+b(p_{t}^{*}-\epsilon_{t}))\cdot b(p_{2}-p_{1})$
    $>$ $\frac{F(\gamma_{t}-a+bp_{t}^{*})}{2}\cdot b(p_{2}-p_{1})$
    $\geq$ $C_{\epsilon}\cdot(p_{2}-p_{1}).$
    Here the third line is because $0<p_{1}<p_{2}$ and therefore $0<F(\gamma_{t}-a+bp_{1})\leq F(\gamma_{t}-a+bp_{2})$.
  7. 7.
    According to part (6), for any $p\in(p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t})$, we have $|r^{\prime}_{t}(p)|=|r^{\prime}_{t}(p)-r^{\prime}_{t}(p_{t}^{*})|\geq C_{\epsilon}\cdot|p-p_{t}^{*}|$. Therefore, we have
    $(r^{\prime}_{t}(p))^{2}\geq C_{\epsilon}^{2}\cdot(p-p_{t}^{*})^{2}\geq\frac{C_{\epsilon}^{2}}{\frac{L_{r}}{2}}\cdot(r_{t}(p_{t}^{*})-r_{t}(p))=\frac{2C_{\epsilon}^{2}}{L_{r}}\cdot(r_{t}(p_{t}^{*})-r_{t}(p)).$ (21)
    Let $C_{v}:=\frac{L_{r}}{2C_{\epsilon}^{2}}$ and the property is proven.
A.2 Proof of Lemma 5.3
Proof.
Recall that $\gamma_{0}:=a_{\max}-b_{\min}p_{\max}+c$. Notice that
$\frac{e_{1,t}-e_{2,t}}{\gamma_{t}-\gamma_{0}}\in[0,\frac{1}{\gamma_{\min}-\gamma_{0}}],e_{3,t}\in[0,1].$ (22)
Also, for any $t=1,2,\ldots,\tau$ in Stage 1, and any $\gamma$ such that $\gamma_{\min}\leq\gamma_{t}$, we have
$\mathbb{E}[\mathds{1}[D_{t}\geq\gamma]]=$ $\mathbb{E}_{N_{t}}[\mathbb{E}_{p_{t}\sim U[0,p_{\max}]}[a-bp_{t}+N_{t}\geq\gamma|N_{t}]]$ (23)
$=$ $\mathbb{E}_{N_{t}}[\mathbb{E}_{p_{t}\sim U[0,p_{\max}]}[p_{t}\leq\frac{a-\gamma+N_{t}}{b}|N_{t}]]$
$=$ $\mathbb{E}_{N_{t}}[\frac{a-\gamma+N_{t}}{b\cdot p_{\max}}]$
$=$ $\frac{a-\gamma}{bp_{\max}}.$
Here the first row is due to Law of Total Expectation, and the last row is due to the zero-mean assumption of $N_{t}$ (see Assumption 3.6). Given this equation, we have:
$\mathbb{E}[\frac{e_{1,t}-e_{2,t}}{\gamma_{t}-\gamma_{0}}]$ $=\frac{\frac{\frac{2\gamma_{t}+2\gamma_{0}}{4}-\frac{1\gamma_{t}+3\gamma_{0}}{4}}{bp_{\max}}}{\gamma_{t}-\gamma_{0}}=\frac{\frac{1}{4}\cdot(\gamma_{t}-\gamma_{0}\cdot\frac{1}{bp_{\max}})}{\gamma_{t}-\gamma_{0}}=\frac{1}{4bp_{\max}}.$ (24)
$\mathbb{E}[e_{3,t}]$ $=\frac{a-\frac{3\gamma_{t}+\gamma_{0}}{4}}{bp_{\max}}.$
According to Hoeffding’s Inequality, we have with $\Pr\geq 1-\eta\delta$:
$|\frac{1}{\tau}\sum_{t=1}^{\tau}\frac{e_{1,t}-e_{2,t}}{\gamma_{t}-\gamma_{0}}-\frac{1}{4bp_{\max}}|\leq\frac{1}{\gamma_{\min}-\gamma_{0}}\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}\cdot\frac{1}{\sqrt{\tau}}.$ (25)
Based on this concentration, we upper bound the estimation error between $b$ and $\hat{b}$ by the end of Stage 1:
$|\hat{b}-b|$ $=|\frac{1}{4p_{\max}\cdot\frac{1}{\tau}\sum_{t=1}^{\tau}\frac{e_{1,t}-e_{2,t}}{\gamma_{t}-\gamma_{0}}}-b|$ (26)
$=|\frac{1}{4p_{\max}(\frac{1}{\tau}\sum_{t=1}^{\tau}\frac{e_{1,t}-e_{2,t}}{\gamma_{t}-\gamma_{0}}-\frac{1}{4bp_{\max}})+\frac{1}{b}}-\frac{1}{\frac{1}{b}}|$
$=|\frac{4p_{\max}\cdot\frac{1}{\tau}\sum_{t=1}^{\tau}\frac{e_{1,t}-e_{2,t}}{\gamma_{t}-\gamma_{0}}\cdot\frac{1}{4bp_{\max}}}{4p_{\max}\cdot\frac{1}{\tau}\sum_{t=1}^{\tau}\frac{e_{1,t}-e_{2,t}}{\gamma_{t}-\gamma_{0}}}|$
$\leq\frac{4p_{\max}\frac{1}{\gamma_{\min}-\gamma_{0}}\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}\cdot\frac{1}{\sqrt{\tau}}}{4p_{\max}(\frac{1}{4bp_{\max}}-\frac{1}{\gamma_{\min}-\gamma_{0}}\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}\cdot\frac{1}{\sqrt{\tau}})\cdot\frac{1}{b}}$
$\leq\frac{\frac{1}{\gamma_{\min}-\gamma_{0}}\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}\cdot\frac{1}{\sqrt{\tau}}}{\frac{1}{8bp_{\max}}\cdot\frac{1}{b}}$
$\leq\frac{4b_{\max}^{2}p_{\max}}{\gamma_{\min}-\gamma_{0}}\cdot\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}\cdot\frac{1}{\sqrt{\tau}}.$
Here the fifth row requires $\frac{1}{8bp_{\max}}\geq\frac{1}{\gamma_{\min}-\gamma_{0}}\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}\cdot\frac{1}{\sqrt{\tau}}$, which further requires $T\geq(\frac{8bp_{\max}}{\gamma_{\min}-\gamma_{0}})^{4}\cdot\frac{1}{4}\log^{2}\frac{2}{\eta\theta}$. According to Assumption 3.8, we know that this inequality holds. Denote $C_{b}:=\frac{4b_{\max}^{2}p_{\max}}{\gamma_{\min}-\gamma_{0}}\cdot\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}$ and we have $|\hat{b}-b|\leq C_{b}\cdot\frac{1}{\sqrt{\tau}}$ with high probability.
Again, according to Hoeffding’s Inequality, we have with $\Pr\geq 1-\eta\delta$:
$|\frac{1}{\tau}\sum_{t=1}^{\tau}e_{3,t}-\frac{a-\frac{3\gamma_{t}+\gamma_{0}}{4}}{bp_{\max}}|\leq\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}\cdot\frac{1}{\sqrt{\tau}}.$ (27)
Hence we have
$|\hat{a}-a|=$ $|\frac{1}{\tau}\sum_{t=1}^{\tau}(\hat{b}p_{\max}e_{3,t}+\frac{3\gamma_{t}+\gamma_{0}}{4})-a|$ (28)
$=$ $|(\hat{b}-b)p_{\max}\cdot\frac{1}{\tau}\sum_{t=1}^{\tau}e_{3,t}+\frac{1}{\tau}\sum_{t=1}^{\tau}(bp_{\max}e_{3,t}-(a-\frac{3\gamma_{t}+\gamma_{0}}{4}))|$
$\leq$ $|\hat{b}-b|p_{\max}\cdot 1+|\frac{1}{\tau}\sum_{t=1}^{\tau}bp_{\max}e_{3,t}-(a-\frac{3\gamma_{t}+\gamma_{0}}{4})|$
$\leq$ $p_{\max}|\hat{b}-b|+bp_{\max}\cdot\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}\cdot\frac{1}{\sqrt{\tau}}$
$=$ $p_{\max}\cdot C_{b}\cdot p_{\max}\frac{1}{\sqrt{\tau}}+bp_{\max}\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}\cdot\frac{1}{\sqrt{\tau}}$
$\leq$ $p_{\max}(C_{b}+b_{\max}\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}})\cdot\frac{1}{\sqrt{\tau}}.$
Denote $C_{a}:=p_{\max}(C_{b}+b_{\max}\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}})$ and the lemma is proven. ∎
A.3 Proof of Lemma 5.4
Proof.
Here we consider the time periods before time $t$, and we use an index $s$ to denote each time period $s=1,2,\ldots,t-1,t$. As a consequence, we have the notations $D_{s},\gamma_{s},k_{s},\mathds{1}_{s}$ corresponding to $D_{t},\gamma_{t},k_{t},\mathds{1}_{t}$ as we defined in Section 3 and Section 4. Also, we denote $N_{k}(t),F_{k}(t)$ and $G_{k}(t)$ as the value of $N_{k},F_{k}$ and $G_{k}$ at the beginning of time period $t$.
From Algorithm 1, we have
$|\hat{r}_{k,t}-r^{\prime}_{t}(p_{k,t})|\leq$ $|G_{k}(t)-(G(c)-G(\gamma_{t}-a+bp_{k,t}))|+p_{k,t}|\hat{b}F_{k}-bF(\gamma_{t}-a+bp_{k,t})|.$ (29)
Notice that $G_{k}(t)=\frac{1}{N_{k}(t)}\cdot\sum_{s=1}^{t-1}\mathds{1}[k_{s}==k]\cdot(D_{s}-\gamma_{s}+c)$. Also, for each $D_{s}$ on the price $p_{k_{s},s}$, we have
$\mathbb{E}[D_{s}|p_{k_{s},s}=\frac{W_{k}-(\gamma_{s}-\hat{a})}{\hat{b}}]=(\gamma_{t}-c+G(c)-G(\gamma_{s}-a+bp_{k_{s},s})).$ (30)
Recall that $W_{k}=2k\Delta$. Hence we have
$\mathbb{E}[G_{k}(t)]=$ $\frac{1}{N_{k}(t)}\sum_{s=1}^{t-1}\mathds{1}[k_{s}==k]\cdot\mathbb{E}[D_{s}-\gamma_{s}+c]$ (31)
$=$ $\frac{1}{N_{k}(t)}\sum_{s=1}^{t-1}\mathds{1}[k_{s}==k](G(c)-G(\gamma_{s}-a+b\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{\hat{b}})).$
Also, due to the fact that $G(x)$ is $1$-Lipschitz, we have
$|G(\gamma_{s}-a+b\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{\hat{b}})|$ (32)
$=$ $|G(\gamma_{s}-a+b\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{b}+b(\frac{1}{\hat{b}}-\frac{1}{b})(k\delta-\gamma_{s}+\hat{a}))-G(2k\Delta)|$
$=$ $|G(\gamma_{s}-a+2k\Delta-\gamma_{s}+\hat{a}+\frac{b-\hat{b}}{\hat{b}}\cdot(2k\Delta-\gamma_{s}+\hat{a}))-G(2k\Delta)|$
$=$ $|G(2k\Delta+(\hat{a}-a)+(b-\hat{b})\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{\hat{b}})-G(2k\Delta)|$
$\leq$ $|2k\Delta+(\hat{a}-a)+(b-\hat{b})\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{\hat{b}}-2k\Delta|$
$\leq$ $|\hat{a}-a|+|b-\hat{b}|\cdot\frac{|2k\Delta-\gamma_{s}+\hat{a}|}{\hat{b}}$
$<$ $(C_{a}+C_{b}\cdot\frac{c+a_{\max}}{b_{\min}})\cdot\frac{1}{\sqrt{\tau}}.$
Therefore, we know that
$|\mathbb{E}[G_{k}(t)]-(G(c)-G(2k\Delta))|\leq$ $\frac{1}{N_{k}(t)}\sum_{s=1}^{t-1}\mathds{1}[k_{s}==k](C_{a}+C_{b}\cdot\frac{c+a_{\max}}{b_{\min}})\cdot\frac{1}{\sqrt{\tau}}$ (33)
$\leq$ $(C_{a}+C_{b}\cdot\frac{c+a_{\max}}{b_{\min}})\cdot\frac{1}{\sqrt{\tau}}.$
Also, since each $D_{s}-\gamma_{s}+c<a_{max}+c$, according to Hoeffding’s Inequality, with $\Pr\geq 1-\eta\delta$ we have
$|G_{k}(t)-(G(c)-G(2k\Delta))|$ (34)
$\leq$ $|G_{k}(t)-\mathbb{E}[G_{k}(t)]|+|\mathbb{E}[G_{k}(t)]-(G(c)-G(2k\Delta))|$
$\leq$ $(a_{\max}+c)\cdot\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}\frac{1}{\sqrt{N_{k}{t}}}+(C_{a}+C_{b}\cdot\frac{c+a_{\max}}{b_{\min}})\cdot\frac{1}{\sqrt{\tau}}.$
On the other hand, since
$\mathbb{E}[F_{k}(t)]=$ $\frac{1}{N_{k}(t)}\sum_{s=1}^{t-1}\mathds{1}[k_{s}==k]\mathbb{E}[\mathds{1}[D_{s}<\gamma_{s}]]$ (35)
$=$ $\frac{1}{N_{k}(t)}\sum_{s=1}^{t-1}\mathds{1}[k_{s}==k]\cdot F(\gamma_{s}-a+b\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{\hat{b}}).$
Similar to Eq. 32, since $F(x)$ is $L_{F}$-Lipschitz, we have
$|F(\gamma_{s}-a+b\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{\hat{b}})-F(2k\Delta)|$ (36)
$=$ $|F(\gamma_{s}-a+b\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{b}+b(\frac{1}{\hat{b}}-\frac{1}{b})(2k\Delta-\gamma_{s}+\hat{a}))-F(2k\Delta)|$
$=$ $|F(\gamma_{s}-a+2k\Delta-\gamma_{s}+\hat{a}+\frac{b-\hat{b}}{\hat{b}}(2k\Delta-\gamma_{s}+\hat{a}))-F(2k\Delta)|$
$=$ $|F(2k\Delta+(\hat{a}-a)+(b-\hat{b})\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{\hat{b}})-F(2k\Delta)|$
$\leq$ $L_{F}\cdot(|\hat{a}-a|+|b-\hat{b}|\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{\hat{b}})$
$<$ $L_{F}\cdot(C_{a}+C_{b}\cdot\frac{c+a_{\max}}{b_{\min}})\cdot\frac{1}{\sqrt{\tau}}.$
Therefore, we have
$|\mathbb{E}[F_{k}(t)-F(2k\Delta)]|\leq L_{F}\cdot(C_{a}+C_{b}\cdot\frac{c+a_{\max}}{b_{\min}})\cdot\frac{1}{\sqrt{\tau}}.$ (37)
Also, since $\mathds{1}_{s}=\mathds{1}[D_{s}<\gamma_{s}]$, according to Hoeffding’s inequality, with $\Pr\geq 1-\eta\delta$ we have
$|F_{k}(t)-F(2k\Delta)|$ (38)
$\leq$ $|F_{k}(t)-\mathbb{E}[F_{k}(t)]|+|\mathbb{E}[F_{k}(t)]-F(2k\Delta)|$
$\leq$ $1\cdot\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}\cdot\frac{1}{N_{k}(t)}+L_{F}(C_{a}+C_{b}\cdot\frac{c+a_{\max}}{b_{\min}})\cdot\frac{1}{\sqrt{\tau}}.$
As a consequence, we have bounded the estimation error of $\hat{r}_{k,t}$ from $r^{\prime}_{t}(p_{k,t})$:
$|r^{\prime}_{t}(p_{k,t})-\hat{r}_{k,t}|$ (39)
$\leq$ $|\gamma_{t}-c+G_{k}(t)-\hat{b}p_{k,t}F_{k}(t)-(\gamma_{t}-c+G(c)-G(2k\Delta)-bp_{k,t}F(2k\Delta))|$
$+|G(2k\Delta)-G(\gamma_{t}-a+b\cdot\frac{2k\Delta-(\gamma_{t}-\hat{a})}{\hat{b}})|+bp_{k,t}|F(2k\Delta)-F(\gamma_{t}-a+b\cdot\frac{2k\Delta-(\gamma_{t}-\hat{a})}{\hat{b}})|$
$\leq$ $|G_{k}(t)-(G(c)-G(2k\Delta))|+|\hat{b}-b|\cdot p_{k,t}F_{k}+bp_{k,t}|F_{k}(t)-F(2k\Delta)|$
$+|G(2k\Delta)-G(\gamma_{t}-a+b\cdot\frac{2k\Delta-(\gamma_{t}-\hat{a})}{\hat{b}})|+b_{\max}p_{\max}\cdot|F(2k\Delta)-F(\gamma_{t}-a+b\cdot\frac{2k\Delta-(\gamma_{t}-\hat{a})}{\hat{b}})|$
$\leq$ $(c+a_{\max})\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}\cdot\frac{1}{\sqrt{N_{k}(t)}}+(C_{a}+C_{b}\cdot\frac{c+a_{\max}}{b_{\min}})\cdot\frac{1}{\sqrt{\tau}}+p_{\max}C_{b}\cdot\frac{1}{\sqrt{\tau}}$
$+b_{\max}p_{\max}(\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}\cdot\frac{1}{\sqrt{N_{k}(t)}}+L_{F}(C_{a}+C_{b}\cdot\frac{c+a_{\max}}{b_{\min}})\cdot\frac{1}{\sqrt{\tau}})$
$+(C_{a}+C_{b}\cdot\frac{c+a_{\max}}{b_{\min}})\cdot\frac{1}{\sqrt{\tau}}+b_{\max}p_{\max}L_{F}(C_{a}+C_{b}\cdot\frac{c+a_{\max}}{b_{\min}})\cdot\frac{1}{\tau}$
=:absent:=:= : $C_{N}\cdot\frac{1}{\sqrt{N_{k}(t)}}+C_{\tau}\cdot\frac{1}{\sqrt{\tau}}=\Delta_{k}.$
Here
$C_{N}$ $:=(c+a_{\max}+b_{\max}p_{\max})\cdot\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}},$ (40)
$C_{\tau}$ $:=2(b_{\max}p_{\max}L_{F}+1)(C_{a}+C_{b}\cdot\frac{c+a_{\max}}{b_{\min}})+p_{\max}C_{b}.$
Finally, we apply the union bound on the probability, and know that Eq. 39 holds with probability $\Pr\geq 1-6\eta\delta$. ∎
A.4 Proof of Lemma 5.5
Proof.
We first prove the lemma under Case 1 when some confidence bound contains 00. Denote $\rho_{t}:=\min\{|r^{\prime}_{t}(p_{t}^{*}-\epsilon_{t})|,|r^{\prime}_{t}(p_{t}^{*}+\epsilon_{t})|\}$, and we know that $\rho_{t}>0$ due to the uniqueness of $p_{t}^{*}$.
Now, let $N_{0}:=\frac{36C_{N}^{2}}{\rho_{t}^{2}}$, where $C_{N}$ is the constant coefficient define in Lemma 5.4. Given this, when $N_{k}(t)\geq N_{0}$, we have
$C_{N}\cdot\frac{1}{\sqrt{N_{k}(t)}}$ $\leq C_{N}\frac{1}{\sqrt{N_{0}}}$ (41)
$=C_{N}\cdot\frac{1}{\frac{6C_{N}}{\rho_{t}}}$
$=C_{N}\cdot\frac{\rho_{t}}{6C_{N}}=\frac{\rho_{t}}{6}.$
Also, since $T$ is assumed as larger than any constant (see Assumption 3.8), we have $T\geq\frac{1296C_{\tau}^{4}}{\rho_{t}^{4}}$, and therefore
$C_{\tau}\cdot\frac{1}{\sqrt{\tau}}$ $=C_{\tau}\frac{1}{T^{1/4}}$ (42)
$\leq C_{\tau}\cdot\frac{1}{\frac{6C_{\tau}}{\rho_{t}}}$
$=C_{N}\cdot\frac{\rho_{t}}{6C_{\tau}}=\frac{\rho_{t}}{6}.$
Given eq. 41 and eq. 42, we know that $\Delta_{k}(t)=C_{N}\cdot\frac{1}{\sqrt{N_{k}(t)}}+C_{\tau}\cdot\frac{1}{\sqrt{\tau}}\leq\frac{\rho_{t}}{6}+\frac{\rho_{t}}{6}=\frac{\rho_{t}}{3}$. Now, if $0\in[\hat{r}_{k,t}-\Delta_{k}(t),\hat{r}_{k,t}+\Delta_{k}(t)]$, we have $|\hat{r}_{k,t}|\leq\Delta_{k}(t)$ and therefore
$|r^{\prime}_{t}(p_{k,t})|\leq|r^{\prime}_{t}(p_{k,t})-\hat{r}_{k,t}|+|\hat{r}_{k,t}-0|\leq\Delta_{k}(t)+\Delta_{k}(t)\leq\frac{\rho_{t}}{3}+\frac{\rho_{t}}{3}=\frac{2\rho_{t}}{3}.$ (43)
Since $r^{\prime}_{t}(p)$ is monotonically non-increasing, any $p\in[0,p_{\max}]$ satisfying $r^{\prime}_{t}(p)<\rho_{t}$ should satisfy $p\in(p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t})$. Therefore, we have $p_{k,t}\in(p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t})$. According to Lemma 5.2 Property (7), we have:
$|r_{t}(p_{t}^{*})-r_{t}(p_{k,t})|\leq$ $C_{v}\cdot(r^{\prime}_{t}(p_{k,t}))^{2}$ (44)
$\leq$ $C_{v}\cdot(2\Delta_{k}(t))^{2}$
$\leq$ $4C_{v}(C_{N}\cdot\frac{1}{\sqrt{N_{k}(t)}}+C_{\tau}\cdot\frac{1}{\sqrt{\tau}})^{2}$
$\leq$ $8C_{v}(C_{N}^{2}\cdot\frac{1}{N_{k}(t)}+C_{\tau}^{2}\cdot\frac{1}{\tau}).$
Let $C_{in}:=8C_{v}\max\{C_{N}^{2},C_{\tau}^{2}\}$ and the first part of Lemma 5.5 holds.
Now let us prove the lemma under Case 2 when no confidence bound contains 0. Formally stated, we have
$(\hat{r}_{k,t}+\Delta_{k}(t))(\hat{r}_{k,t}-\Delta_{k}(t))>0,\forall k=-M,-M+1,\ldots,M-1,M.$ (45)
Denote $\theta_{t}:=\inf_{k}\min\{|\hat{r}_{k,t}+\Delta_{k}(t)|,|\hat{r}_{k,t}-\Delta_{k}(t)|\}$, and we know that $\min\{|\hat{r}_{k_{t},t}+\Delta_{k}(t)|,|\hat{r}_{k_{t},t}-\Delta_{k}(t)|\}\theta_{t}>0$, where $k_{t}$ is the $k$ such that $p_{k,t}$ is proposed at time $t$. Therefore, we have $|\hat{r}_{k,t}+\Delta_{k}(t)|\geq\theta_{t}$ and $|\hat{r}_{k,t}-\Delta_{k}(t)|\geq\theta_{t},\forall k$. According to the prerequisite of Lemma 5.5 Part (2), there exists $k_{0}$ such that
[…]
Also, since $r^{\prime}_{t}(p)$ is $L_{r}$-Lipschitz, we have
$|r^{\prime}_{t}(p_{k_{0},t})-r^{\prime}_{t}(p_{k_{0}+1,t})|$ $\leq L_{r}(p_{k_{0}+1,t}-p_{k_{0},t})=L_{r}\frac{2\Delta}{\hat{b}}$ (47)
$\Rightarrow\quad 2L_{r}\frac{\Delta}{\hat{b}}$ $\geq|r^{\prime}_{t}(p_{k_{0},t})-r^{\prime}_{t}(p_{k_{0}+1,t})|\geq 2\theta_{t}$
$\Rightarrow\quad\theta_{t}$ $\leq\frac{L_{r}}{\hat{b}}\cdot\Delta\leq\frac{L_{r}}{b_{\min}}(C_{a}+C_{b}p_{\max})\cdot\frac{1}{T^{1/4}}.$
As $T$ is sufficiently large, we have $\frac{L_{r}}{b_{\min}}(C_{a}+C_{b}p_{\max})\cdot\frac{1}{T^{1/4}}\leq\frac{1}{6}\cdot\rho_{t}$ where $\rho_{t}:=\min\{|r^{\prime}_{t}(p_{t}^{*}-\epsilon_{t})|,|r^{\prime}_{t}(p_{t}^{*}+\epsilon_{t})|\}$. Let $N_{1}:=\frac{144C_{N}^{2}}{\rho_{t}^{2}}$. Similar to Eq. 41 and Eq. 42, we have
$C_{N}\cdot\frac{1}{\sqrt{N_{k_{t}}(t)}}\leq$ $\frac{1}{12}\rho_{t}$ (48)
$C_{\tau}\cdot\frac{1}{\sqrt{\tau}}\leq$ $\frac{1}{12}\rho_{t}.$
Hence, we have
$|r^{\prime}_{t}(p_{k_{t},t})|\leq 2\Delta_{k_{t}}(t)+\theta_{t}\leq 2\cdot(\frac{1}{12}\rho_{t}+\frac{1}{12}\rho_{t})+\frac{\rho_{t}}{6}=\frac{\rho_{t}}{2}.$ (49)
Since $r^{\prime}_{t}(p)$ is monotonically non-increasing, we know that $p_{k_{t},t}\in(p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t})$ similar to the analysis in Case (1), and again we have $|r_{t}(p_{t}^{*})-r_{t}(p_{k_{t},t})|\leq 8C_{v}(C_{N}^{2}\cdot\frac{1}{N_{k_{t}}(t)}+C_{\tau}^{2}\cdot\frac{1}{\tau})$. Now we complete the proof of Lemma 5.5 on both circumstances. ∎
A.5 Proof of Lemma 5.6
Proof.
When $\gamma_{t}>\frac{\hat{a}+C_{a}\cdot\frac{1}{\sqrt{\tau}}}{2}+c>\frac{a}{2}+c$, we know that the demand at $p=\frac{a}{2b}$ and its neighborhood is not censored. As a result, $\frac{a}{2b}$ is still a local optimal (and therefore global optimal due to the unimodality) of $r_{t}(p)$. According to Lemma 5.3, we have:
$r_{t}(\frac{a}{2b})-r_{t}(\frac{\hat{a}}{2\hat{b}})$ $\leq C_{s}(\frac{a}{2b}-\frac{\hat{a}}{2\hat{b}})^{2}$ (50)
$=C_{s}((\frac{a}{2b}-\frac{a}{2\hat{b}})+(\frac{a}{2\hat{b}}-\frac{\hat{a}}{2\hat{b}}))^{2}$
$\leq C_{s}(\frac{a|\hat{b}-b|}{2b\hat{b}}+\frac{a-\hat{a}}{2\hat{b}})^{2}$
$\leq 2C_{s}(\frac{a_{\max}^{2}}{4b_{\min}^{4}}(\hat{b}-b)^{2}+\frac{1}{4b_{\min}^{2}}(a-\hat{a})^{2})$
$\leq\frac{2C_{s}(a_{\max}^{2}C_{b}^{2}+b_{\min}^{2}C_{a}^{2})}{4b_{\min}^{4}}\cdot\frac{1}{\tau}.$
When $\hat{r}_{k,t}-\Delta_{k}(t)>0,\forall k=-M,-M+1,\ldots,M-1,M$, we know that $r^{\prime}_{t}(\frac{a+c-\gamma_{t}}{b})>0$ according to Lemma 5.4. Since $d_{t}(\frac{a+c-\gamma_{t}}{b})=a-b\cdot\frac{a+c-\gamma_{t}}{b}+N_{t}=\gamma_{t}-c+N_{t}\leq\gamma_{t}$ is not censored, we know that the optimal price $p_{t}^{*}$ satisfies $p_{t}^{*}>\frac{a+c-\gamma_{t}}{b}$ (since $r^{\prime}_{t}(p_{t}^{*})=0<r^{\prime}_{t}(\frac{a+c-\gamma_{t}}{b})$) and therefore its demand is not censored. Therefore, the optimal price $p_{t}^{*}=\frac{a}{2b}$ and we have its regret bounded by Eq. 50 identically. Let $C_{non}:=\frac{2C_{s}(a_{\max}^{2}C_{b}^{2}+b_{\min}^{2}C_{a}^{2})}{4b_{\min}^{4}}$ and we have proven both cases. ∎